# Growth problems for representations of finite monoids

### v1

@@ 1. Introduction
[…]
@@
The questions (1)(3) above are a special case of growth problems, which following [LTV24] we may define for any additive Krull–Schmidt monoidal category. The growth problems and related questions for various categories have been recently studied in the papers [CEO24, CEOT24, He25, KST24, LTV23, LTV24, LPRS24, Lar24]. In particular, if our monoid $M$ is a group, then the situation is well-understood from [He25, Theorem 1], which gives an expression for the asymptotic formula $a(n)$ in terms of the (Brauer) character table. A main goal of this paper is to generalize this result to an arbitrary finite monoid.
@@
1B. A conjecture and evidence
Let $G$ denote the group of units of $M$. Let $Z_{V}(G)$ denote the set of elements $g$ in $G$ which act as scalars on $V$, and denote the corresponding scalars by $\omega_{V}(g)$. Let $g_{t},1\leq t\leq N$ be a complete set of representatives for the $p$-regular conjugacy classes of $G$ (if $p=0$, these are all the conjugacy classes). If $V$ is a $kM$-module, recall from [LTV24, Section 2] that the corresponding fusion graph $\Gamma$ is the (oriented and weighted) graph whose vertices are indecomposable $kM$-modules which are summands of $V^{\otimes n}$, for some $n\geq 0$, and that there is an edge of weight $m$ from the vertex $V_{i}$ to the vertex $V_{j}$ if $V_{i}$ occurs $m$ times in the direct sum decomposition of $V\otimes V_{j}$. The corresponding (potentially countably infinite) adjacency matrix is called the action matrix. If $V$ is a $kM$-module, let $\operatorname{Res}_{G}(V)$ denote the $kG$-module that comes from restricting the action on $V$. Let $\lambda^{\mathrm{sec}}$ denote any second largest eigenvalue (in terms of modulus) of the action matrix of $V$. We use the usual capital $O$ notation. We conjecture that the asymptotic growth rate $a(n)$ is the same as that of $\operatorname{Res}_{G}(V)$.
[…]
@@
    (1B.1) $a(n)=\frac{1}{\lvert G\rvert}\sum_{\begin{subarray}{c}1\leq t\leq N\\ g_{t}\in Z_{V}(G)\end{subarray}}S_{t}\big{(}\omega_{V}(g_{t})\big{)}^{n}\cdot(\dim V)^{n},$
    where $S_{t}$ is the sum over entries of the column corresponding to $g_{t}^{-1}$ in the (irreducible) Brauer character table. In other words, $V$ has the same asymptotic growth rate as the $kG$-module $\operatorname{Res}_{G}(V)$, c.f. [He25, Theorem 1].
  2. (b)
@@
As evidence for Section 1B, over fixed finite fields $k$ (e.g. $\mathbb{F}_{11}$) we computed the growth rates of all projective indecomposable $kM$-modules, where $M$ is either a monoid of order $\leq 7$ (34129 of these) or is the direct product $T_{3}\times L$ where $T_{3}$ is the full transformation monoid on 3 elements, and $L$ is a monoid of order $\leq 5$ (273 of these). We were not able to find a counterexample to the formula in Equation 1B.1. To compute the growth rates, we used GAP’s Smallsemi package to obtain the multiplication table for the monoids of small order, and from these we constructed the monoid algebras in Magma as matrix algebras. Finally, we computed the action matrix as in the appendix of [He25]. We performed the calculations over finite fields as Magma’s Meataxe algorithm for decomposing modules is much slower over $\mathbb{Q}$.
[…]
The code used in this paper is available on [HT25]. $\Diamond$
1C. Main results
In Section 3A, we prove Section 1B under the hypothesis that some projective $kG$-module is injective as a $kM$-module. While the hypothesis is somewhat restrictive, it is satisfied by important classes of monoids such as the full transformation monoids $T_{n}$ (in characteristic 0), the monoid of $2\times 2$ matrices $\operatorname{M}(2,q)$ (over defining characteristic), and all monoids with semisimple monoid algebra. We compute asymptotic formulas for these examples in Section 4. In addition, we obtain general (exact and asymptotic) formulas for $l(n)$ when the characteristic of $k$ does not divide the order of any of $M$’s maximal subgroups. As $b(n)=l(n)$ when $kM$ is semisimple, we obtain a formula for $b(n)$ in this case, which generalizes [He25, Theorem 3].
Acknowledgments. We would like to express our gratitude to Volodymyr (Walter) Mazorchuk for insightful email exchanges on the topic of monoid representations. In particular, the observation discussed in Section 4C is credited to Walter.
@@ 2. Growth rate in general
The upper bound is immediate, and it remains to justify the lower bound. Moreover, $b_{n}\leq l_{n}$, so it remains to justify the lower bound for $b_{n}$.
@@
Now, any finite semigroup $S$ with #$\#S=m$ can be realized in $\operatorname{M}(m,\mathbb{F}_{2})$ (via the regular representation over $\mathbb{F}_{2}$), so we get a bound of $b_{n}$ from below by $b_{n}$ for $\operatorname{M}(m,\mathbb{F}_{2})$ which is the same as for $\operatorname{GL}(m,\mathbb{F}_{2})$, by the above. The desired lower bound now follows from [COT24, Proposition 2.2]. ∎
@@ 3. Main results
3A. Counting summands
As before, if $V$ is a $kM$-module let $\operatorname{Res}_{G}(V)$ denote the $kG$-module that comes from restriction of $V$. If $W$ is a $kG$-module, let $\operatorname{Ind}_{G}(W)$ denote the induced $kM$-module on which elements in $M\setminus G$ act as 0. We write $P(W)$ for the projective cover of $W$. We note that induction and restriction give inverse equivalences between the category of $kG$-modules and the full subcategory of $kM$-modules with all nonunits acting as 0. Let $\Gamma^{G}$ denote the subgraph in the fusion graph $\Gamma$ for $V$ induced by the vertices $\operatorname{Ind}_{G}(P(W))$ where $W$ is an irreducible $kG$-module.
@@
(3A.1) $V\otimes\operatorname{Ind}_{G}(P(W))=\operatorname{Ind}_{G}(\operatorname{Res}_{G}(V)\otimes P(W)).$
The following theorem establishes Section 1B under the assumption that some module in $\Gamma^{G}$ is injective.
[…]
@@
then the statements of Section 1B hold.
@@
Let $\Gamma^{G}$ be the subgraph of $\Gamma$ as above, then it is strongly connected since $\operatorname{Res}_{G}(V)$ is faithful by assumption i), and the projective cell in the fusion graph of a faithful $kG$-module is strongly-connected (c.f. the proof of [LTV24, Proposition 4.22]).We will first show that the growth problem for $V$ is sustainably positive recurrent in the sense of [LTV24, Definition 5.5], with $\Gamma^{G}$ being the unique final basic class (FBC).
@@
Next, we claim that under the assumptions of the theorem any vertex $Z$ in $\Gamma$ has a path to $\Gamma^{G}$. For $n\geq 0$, the set $V_{n}\subseteq V^{\otimes n}$ of all vectors $v\in V^{\otimes n}$ on which $M\setminus G$ acts as zero is a submodule. To see this, note that $V_{n}$ is clearly a vector subspace (it contains 00, so is nonempty), and if $v\in V_{n}$, for all $m\in M\setminus G$ and $n\in M$ we have $m(nv)=0$ because $mn\in M\setminus G$, and so $nv\in V_{n}$. By assumption i), $V_{n}$ is nonzero for $n$ large enough, say $n=|M\setminus G|$, since for any $m\in M\setminus G$ acts non-invertibly and we can find $v_{m}\in V$ such that $m\cdot v_{m}=0$, and then the basic tensor over all $v_{m}$’s is in $V_{n}$. Thus, for some $n$ large enough, $V^{\otimes n}$ contains a nonzero submodule $V^{\prime}=V_{n}$ on which all non-units act as zero. Because $\operatorname{Res}_{G}(V)$ is faithful, any projective indecomposable $kG$-module $P$ appears in some tensor power of $\operatorname{Res}_{G}(V)$ (see [BK72, Theorem 2]), say $(\operatorname{Res}_{G}(V))^{\otimes k}$. Now $\operatorname{Res}_{G}(V^{\prime}\otimes V^{\otimes k})=\operatorname{Res}_{G}(V^{\prime})\otimes(\operatorname{Res}_{G}(V))^{\otimes k}$ has a projective indecomposable summand, and so $V^{\prime}\otimes V^{\otimes k}\subseteq V^{\otimes(k+n)}$ has a summand of the form $\operatorname{Ind}_{G}(P(W))$ for some irreducible $kG$-module $W$ (all nonunits act as zero on $V^{\prime}\otimes V^{\otimes k}$, so we recover the same module by restriction followed by induction). By assumption ii), we may take $\operatorname{Ind}_{G}(P(W))$ to be injective, and so it is a summand of $V^{\otimes(k+n)}$. This shows that the trivial module $k$ has a path to $\Gamma^{G}$; since $Z=Z\otimes k$ and $\Gamma^{G}$ is final in $\Gamma$, in fact any indecomposable $kM$-module $Z$ has a path to $\Gamma^{G}$.
Now it follows, by the same proof as [LTV24, Proposition 5.7], that the growth problem for $V$ is sustainably positive recurrent with the FBC also final in $\Gamma$. Statements (2) and (3) of Section 1B now follow from [LTV24, Theorem 5.10]. The first statement follows by the same proof as [He25, Theorem 1], after adapting the statement about left eigenvectors in [He25, Lemma 7] by using (Brauer) characters of monoids rather than characters of groups (the right eigenvectors are the same). For a definition of the Brauer characters of monoids we refer to [Ste23, §5]. ∎
[…]
  1. (a)
    When $kM$ is semisimple, condition i) in Section 3A is equivalent to the requirement that no generalized conjugacy class in $M$ apart from that of the identity has character value $\dim V$. Condition ii) is automatically satisfied in this case. We will be able to say much more if $kM$ semisimple, see Section 3B below.
  2. (b)
@@
  3. (c)
    We require condition ii) so that, assuming also condition i) holds, $V^{\otimes n}$ will contain a summand in $\Gamma^{G}$ for $n$ large enough. We suspect that this is always true, which would imply Section 1B, but we do not know how to prove it in general.
See also Section 4C below. $\Diamond$
3B. Counting lengths
Let $V$ be a $kM$-module. To compute $l(n)$ we count the multiplicities now in the Grothendieck ring $G_{0}(kM)$ defined via short exact sequences, which coincides with the additive Grothendieck ring if $kM$ is semisimple (in this case $l(n)=b(n)$).
@@
If $V$ is a $kM$-module, let $N(V)$ denote the analogue of action matrix for $l(n)$. That is, with respect to the basis of isomorphism classes of irreducible $kM$-modules $\textup{Irr}_{k}(M)=\{S_{1},\dots,S_{z}\}$, the $(i,j)$-th entry of $N(V)$ is the composition multiplicity $[V\otimes S_{j}:S_{i}].$ A right eigenvector of $N(e_{i}V)$ may be identified with an element $\mathbf{x}$ in the Grothendieck ring $G_{0}(kG_{e_{i}})$, considered as a vector with respect to the basis $\textup{Irr}_{k}(G_{e_{i}})$. By naturally identifying $G_{0}(kG_{e_{i}})$ with a subset of $\prod_{i=1}^{m}G_{0}(kG_{e_{i}})$, x can be identified with a longer vector $\tilde{\textbf{x}}$ with respect to the basis $\bigcup_{i=1}^{m}\textup{Irr}_{k}(G_{e_{i}})$ which is supported on that subset.
[…]
@@
First we prove a). For $1\leq i\leq m$, let $G_{0}(kM)_{i}$ denote the inverse image of $G_{0}(kG_{e_{i}})$ (identified as a subset of $\prod_{j=1}^{n}G_{0}(kG_{e_{j}})$) under the restriction map. That is, it is the ideal in $G_{0}(kM)$ consisting of elements $[W]$ such that $[e_{j}W]=0$ for $j\neq i$. We have the commutative diagram
@@
where $1\leq j\leq z$ and $\chi_{V}$ and $\chi_{j}$ denote the characters for $V$ and $S_{j}$ respectively, and $m_{1},\dots,m_{z}$ are representatives for the generalized conjugacy classes of $M$. Taking the $i$-th entry, we get: ($i$-th column of $X(M))$cdot ($j$-th column of $M$) $=\chi_{V}(m_{i})\chi_{j}(m_{i})$, but this just says that the $i$-th column of $X(M)$ is left eigenvector of $N(V)$ with eigenvalue $\chi_{V}(m_{i})$. ∎
For $1\leq i\leq m$, let $C_{i,1},\dots,C_{i,k_{i}}$ denote the conjugacy classes of $G_{e_{i}}$, and pick a complete set of representatives $g_{i,1},\dots,g_{i,k_{i}}$ for these classes. Let $V$ be a $kM$-module with character $\chi_{V}$. By Section 3B, when $p\nmid\lvert G_{e_{i}}\rvert,1\leq i\leq m$ each conjugacy class $C_{i,j}$ is canonically associated with a pair of left and right eigenvectors with eigenvalue $\chi_{V}(g_{i,j})$. We will refer to these as the $(i,j)$-th row of $(X(M))^{T}$ and the $(i,j)$-th column of $L^{-1}\overline{X}$, respectively. If $G_{e_{i}}$ is a maximal subgroup of $M$, extending notation from before we let $Z_{V}(G_{e_{i}})$ denote the elements $g\in G_{e_{i}}$ which act as scalars on $V$, and let $\omega_{V}(g)$ denote the corresponding scalar. We will write $k(n)$ for an asymptotic expression of $l(n)$.
[…]
@@
    (3B.1) $l(n)=\sum_{i=1}^{m}\frac{1}{\lvert G_{e_{i}}\rvert}\sum_{j=1}^{k_{i}}\lvert C_{i,j}\rvert S_{i,j}\big{(}\chi(g_{i,j})\big{)}^{n}$
  2. (b)
@@
    (3B.2) $l(n)\sim k(n):=\sum_{i=1}^{m}\frac{(\dim V)^{n}}{\lvert G_{e_{i}}\rvert}\sum_{j:g_{i,j}\in Z_{V}(G_{e_{i}})}\lvert C_{i,j}\rvert S_{i,j}\big{(}\omega_{V}(g_{i,j})\big{)}^{n}.$
  3. (c)
@@
For the eigenvalue $\chi(g_{i,j})$, choose the left eigenvector $w_{ij}^{T}$ to be the $(i,j)$-th row of $(X(M))^{T}=X^{T}L$, and choose the right eigenvector $v_{i,j}$ to be the $(i,j)$-th column of $L^{-1}\overline{X}$, normalized by a factor of $\lvert C_{i,j}\rvert/\lvert G_{e_{i}}\rvert$. From eigendecomposition we get
@@
[…]
If $M=G$ is a group, then if $p\nmid\lvert G\rvert$ Section 3B gives
@@
where for $1\leq i\leq N$, $t_{i}$ is a representative for the $i$-th conjugacy class $C_{i}$ of $G$, and $S_{t}=\overline{S_{t}}$ is the sum over entries in column of the character table corresponding to $g_{t}$. Thus Section 3B generalizes the statement in [He25, Theorem 3]. $\Diamond$
[…]
As we observed in the beginning of the section, if $kM$ is semisimple then $b(n)=l(n)$, so Section 3B gives a refinement of the result of Section 3A in this case. In particular, if no element other than $1$ act as identity on $M$, then the second statement of Section 3B recovers the formula of $a(n)$ in Section 3A in this case: this is because the submatrix of $L$ corresponding to the group of units $G$ is the identity matrix, so $S_{i,j}$ is exactly a column of the character table of $G$. Of course, if $kM$ is semisimple then we must have $p\nmid\lvert G_{e_{i}}\rvert$ for all $1\leq i\leq m$, so Section 3B applies (see e.g.[Ste16b, Theorem 5.19]). $\Diamond$

### v2

@@ 1.1. Growth problems
$b(n)=b^{M,V}(n)=\text{\# $M$-indecomposable summands in}\ V^{\otimes n}\ \text{(counted with multiplicity)}.$
@@
The questions (1)(3) above are a special case of growth problems, which following [LTV24] we may define for any additive Krull–Schmidt monoidal category. The growth problems and related questions for various categories have been recently studied in the papers [CEO24, CEOT25, He25, KST24, LTV23, LTV24, LPRS24, Lar24]. In particular, if our monoid $M$ is a group, then the situation is well-understood from [He25, Theorem 1], which gives an expression for the asymptotic formula $a(n)$ in terms of the (Brauer) character table. A main goal of this paper is to generalize this result to an arbitrary finite monoid.
@@
1.2. A conjecture and evidence
Let $G$ denote the group of units of $M$. Let $Z_{V}(G)$ denote the set of elements $g$ in $G$ which act as scalars on $V$, and denote the corresponding scalars by $\omega_{V}(g)$. Let $g_{t},1\leq t\leq N$ be a complete set of representatives for the $p$-regular conjugacy classes of $G$ (if $p=0$, these are all the conjugacy classes). If $V$ is a $kM$-module, recall from [LTV24, Section 2] that the corresponding fusion graph $\Gamma$ is the (oriented and weighted) graph whose vertices are indecomposable $kM$-modules which are summands of $V^{\otimes n}$, for some $n\geq 0$, and that there is an edge of weight $m$ from the vertex $V_{i}$ to the vertex $V_{j}$ if $V_{i}$ occurs $m$ times in the direct sum decomposition of $V\otimes V_{j}$. The corresponding (potentially countably infinite) adjacency matrix is called the action matrix. If $V$ is a $kM$-module, let $\operatorname{Res}_{G}(V)$ denote the $kG$-module that comes from restricting the action on $V$. Let $\lambda^{\mathrm{sec}}$ denote any second largest eigenvalue (in terms of modulus) of the action matrix of $V$. We use the usual capital $O$ notation. We conjecture that the asymptotic growth rate $a(n)$ is the same as that of $\operatorname{Res}_{G}(V)$. We note that if $M=G$ is a group, Conjecture 1 recovers the statement of [He25, Theorem 1].
Conjecture 1.
@@ 1.2. A conjecture and evidence
    (1) $a(n)=\frac{1}{\lvert G\rvert}\sum_{\begin{subarray}{c}1\leq t\leq N\\ g_{t}\in Z_{V}(G)\end{subarray}}S_{t}\big(\omega_{V}(g_{t})\big)^{n}\cdot(\dim V)^{n},$
    where $S_{t}$ is the sum over entries of the column corresponding to $g_{t}^{-1}$ in the (irreducible) Brauer character table. In other words, $V$ has the same asymptotic growth rate as the $kG$-module $\operatorname{Res}_{G}(V)$, cf. [He25, Theorem 1].
  2. 2.
@@
The reason we expect Conjecture 1 to be true is the following: The collection $\mathcal{G}$ of indecomposable $kM$-modules which are $kG$-modules (that is, $M\setminus G$ acts as 0) forms a tensor ideal in the sense that if $K\in\mathcal{G}$ and $V$ is any $kM$-module, then $V\otimes K$ is a direct sum of modules in $\mathcal{G}$. By the technical machinery introduced in [LTV24], our conjecture that $\mathcal{G}$ determine the asymptotic growth rate is roughly equivalent to the conjecture that every vertex in the fusion graph has a path to $\mathcal{G}$, which we suspect is true based on experimental evidence.
As evidence for Conjecture 1, over fixed finite fields $k$ (e.g. $\mathbb{F}_{11}$) we computed the growth rates of all projective indecomposable $kM$-modules, where $M$ is either a monoid of order $\leq 7$ (34129 of these) or is the direct product $T_{3}\times L$ where $T_{3}$ is the full transformation monoid on 3 elements, and $L$ is a monoid of order $\leq 5$ (273 of these). We were not able to find a counterexample to the formula in Equation 1. To compute the growth rates, we used GAP’s Smallsemi package to obtain the multiplication table for the monoids of small order, and from these we constructed the monoid algebras in Magma as matrix algebras. Finally, we computed the action matrix as in the appendix of [He25]. We performed the calculations over finite fields as Magma’s Meataxe algorithm for decomposing modules is much slower over $\mathbb{Q}$.
Remark 2.
The code used in this paper is available on the GitHub repository [HT25]. $\Diamond$
1.3. Main results
In Theorem 1, we prove Conjecture 1 under the hypothesis that some projective $kG$-module is injective as a $kM$-module. While the hypothesis is somewhat restrictive, it is satisfied by important classes of monoids such as the full transformation monoids $T_{m}$ (in characteristic 0), the monoid of $2\times 2$ matrices $\operatorname{M}(2,q)$ (over defining characteristic), and all monoids with semisimple monoid algebra. We compute asymptotic formulas for these examples in Section 4. In addition, we obtain general (exact and asymptotic) formulas for $l(n)$ when the characteristic of $k$ does not divide the order of any of $M$’s maximal subgroups. As $b(n)=l(n)$ when $kM$ is semisimple, we obtain a formula for $b(n)$ in this case, which generalizes [He25, Theorem 3].
Acknowledgments. We would like to express our gratitude to Volodymyr (Walter) Mazorchuk for insightful email exchanges on the topic of monoid representations. In particular, the observation discussed in Section 4.3 is credited to Walter. We would also like to thank the referee for giving the paper a careful reading and providing many helpful comments.
@@ 2. Growth rate in general
The upper bound is immediate, and it remains to justify the lower bound. Moreover, $b(n)\leq l(n)$, so it remains to justify the lower bound for $b(n)$.
@@
Now, any finite semigroup $S$ with #S=m\#S=m can be realized in $\operatorname{M}(m,\mathbb{F}_{2})$ (via the regular representation over $\mathbb{F}_{2}$), so we get a bound of $b(n)$ from below by $b(n)$ for $\operatorname{M}(m,\mathbb{F}_{2})$ which is the same as for $\operatorname{GL}(m,\mathbb{F}_{2})$, by the above. The desired lower bound now follows from [COT24, Proposition 2.2]. ∎
@@ 3. Main results
3.1. Counting summands
This section focuses on the study of $b(n)$. As before, if $V$ is a $kM$-module let $\operatorname{Res}_{G}(V)$ denote the $kG$-module that comes from restriction of $V$. If $W$ is a $kG$-module, let $\operatorname{Ind}_{G}(W)$ denote the induced $kM$-module on which elements in $M\setminus G$ act as 0. We write $P(W)$ for the projective cover of $W$. We note that induction and restriction give inverse equivalences between the category of $kG$-modules and the full subcategory of $kM$-modules with all nonunits acting as 0. Let $\Gamma^{G}$ denote the subgraph in the fusion graph $\Gamma$ for $V$ induced by the vertices $\operatorname{Ind}_{G}(P(W))$ where $W$ is an irreducible $kG$-module.
@@ 3.1. Counting summands
(2) $V\otimes\operatorname{Ind}_{G}(P(W))=\operatorname{Ind}_{G}(\operatorname{Res}_{G}(V)\otimes P(W)).$
Recall from [LTV24, §4] that a final basic class (FBC) of a fusion graph is a strongly-connected component whose PF dimension (= spectral radius of the adjacency submatrix in the case when it is finite, the only case we need) is maximal among all strongly-connected components, and moreover has no path to any other strongly-connected component with maximal PF dimension. If there is no path leaving the FBC at all, we say it is final in the whole graph. The following theorem establishes Conjecture 1 under the assumption that some module in $\Gamma^{G}$ is injective.
Theorem 1.
@@
then the statements of Conjecture 1 hold.
@@
Let $\Gamma^{G}$ be the subgraph of $\Gamma$ as above, then it is strongly connected since $\operatorname{Res}_{G}(V)$ is faithful by assumption i), and the projective cell in the fusion graph of a faithful $kG$-module is strongly-connected (cf. the proof of [LTV24, Proposition 4.22]). We will first show that the growth problem for $V$ is sustainably positive recurrent in the sense of [LTV24, Definition 5.5], with $\Gamma^{G}$ being the unique FBC. This allows us to invoke [LTV24, Theorem 5.10] from which the statements of the theorem easily follow.
@@
Next, we claim that under the assumptions of the theorem any vertex $Z$ in $\Gamma$ has a path to $\Gamma^{G}$. For $n\geq 0$, the set $V_{n}\subseteq V^{\otimes n}$ of all vectors $v\in V^{\otimes n}$ on which $M\setminus G$ acts as zero is a submodule. To see this, note that $V_{n}$ is clearly a vector subspace (it contains 0, so is nonempty), and if $v\in V_{n}$, for all $m\in M\setminus G$ and $n\in M$ we have $m(nv)=0$ because $mn\in M\setminus G$, and so $nv\in V_{n}$. By assumption i), $V_{n}$ is nonzero for $n$ large enough, say $n=|M\setminus G|$, since for any $m\in M\setminus G$ acts non-invertibly and we can find $v_{m}\in V$ such that $m\cdot v_{m}=0$, and then the basic tensor over all $v_{m}$’s is in $V_{n}$. Thus, for some $n$ large enough, $V^{\otimes n}$ contains a nonzero submodule $V^{\prime}=V_{n}$ on which all non-units act as zero. Because $\operatorname{Res}_{G}(V)$ is faithful, any projective indecomposable $kG$-module $P$ appears in some tensor power of $\operatorname{Res}_{G}(V)$ (see [BK72, Theorem 2]), say $(\operatorname{Res}_{G}(V))^{\otimes k}$. Now $\operatorname{Res}_{G}(V^{\prime}\otimes V^{\otimes k})=\operatorname{Res}_{G}(V^{\prime})\otimes(\operatorname{Res}_{G}(V))^{\otimes k}$ has a projective indecomposable summand, and so $V^{\prime}\otimes V^{\otimes k}\subseteq V^{\otimes(k+n)}$ has a summand of the form $\operatorname{Ind}_{G}(P(W))$ for some irreducible $kG$-module $W$ (all nonunits act as zero on $V^{\prime}\otimes V^{\otimes k}$, so we recover the same module by restriction followed by induction). By assumption ii), we may take $\operatorname{Ind}_{G}(P(W))$ to be injective, and so it is a summand of $V^{\otimes(k+n)}$. This shows that the trivial module $k$ has a path to $\Gamma^{G}$; since $Z\cong Z\otimes k$ and $\Gamma^{G}$ is final in $\Gamma$, in fact any indecomposable $kM$-module $Z$ has a path to $\Gamma^{G}$ (we know for $N$ large enough, $k\otimes V^{\otimes N}$ has a summand in $\Gamma^{G}$, so $Z\otimes k\otimes V^{\otimes N}$ does too).
Now it follows, by the same proof as [LTV24, Proposition 5.7], that the growth problem for $V$ is sustainably positive recurrent with the FBC also final in $\Gamma$. Statements (2) and (3) of Conjecture 1 now follow from [LTV24, Theorem 5.10]. The first statement follows by the same proof as [He25, Theorem 1], after adapting the statement about left eigenvectors in [He25, Lemma 7] by using (Brauer) characters of monoids rather than characters of groups (the right eigenvectors are the same). For a definition of the Brauer characters of monoids we refer to [Ste23, §5]. ∎
Remark 3.
  1. 1.
    When kMkM is semisimple, condition i) in Theorem 1 is equivalent to the requirement that no generalized conjugacy class in $M$ apart from that of the identity has character value $\dim V$. Condition ii) is automatically satisfied in this case. We will be able to say much more if kMkM semisimple, see Theorem 2 below.
  2. 2.
@@
  3. 3.
    We require condition ii) so that, assuming also condition i) holds, $V^{\otimes n}$ will contain a summand in $\Gamma^{G}$ for $n$ large enough. We suspect that this is always true, which would imply Conjecture 1, but we do not know how to prove it in general. See also Section 4.3 below.
$\Diamond$
3.2. Counting lengths
We now turn to study the related statistic of $l(n)$. Let $V$ be a $kM$-module. To compute $l(n)$ we count the multiplicities now in the Grothendieck ring $G_{0}(kM)$ defined via short exact sequences, which coincides with the additive Grothendieck ring if $kM$ is semisimple (in this case $l(n)=b(n)$).
@@ 3.2. Counting lengths
If $V$ is a $kM$-module, let $N(V)$ denote the analogue of action matrix for $l(n)$. That is, with respect to the basis of isomorphism classes of irreducible $kM$-modules $\textup{Irr}_{k}(M)=\{S_{1},\dots,S_{z}\}$, the $(i,j)$-th entry of $N(V)$ is the composition multiplicity $V\otimes S_{j}:S_{i}].$ A right eigenvector of $N(e_{i}V)$ may be identified with an element $\mathbf{x}$ in the Grothendieck ring $G_{0}(kG_{e_{i}})$, considered as a vector with respect to the basis $\textup{Irr}_{k}(G_{e_{i}})$. By naturally identifying $G_{0}(kG_{e_{i}})$ with a subset of $\prod_{i=1}^{m}G_{0}(kG_{e_{i}})$, x can be identified with a longer vector $\tilde{\textbf{x}}$ with respect to the basis $\bigcup_{i=1}^{m}\textup{Irr}_{k}(G_{e_{i}})$ which is supported on that subset.
Lemma 1.
@@
First we prove a). For $1\leq i\leq m$, let $G_{0}(kM)_{i}$ denote the inverse image of $G_{0}(kG_{e_{i}})$ (identified as a subset of $\prod_{j=1}^{n}G_{0}(kG_{e_{j}})$) under the restriction map. That is, it is the ideal in $G_{0}(kM)$ consisting of elements $[W]$ such that $e_{j}W]=0$ for $j\neq i$. We have the commutative diagram
@@
where $1\leq j\leq z$ and $\chi_{V}$ and $\chi_{j}$ denote the characters for $V$ and $S_{j}$ respectively, and $m_{1},\dots,m_{z}$ are representatives for the generalized conjugacy classes of $M$. Taking the $i$-th entry, we get: ($i$-th column of $X(M))$ $\cdot$ ($j$-th column of $M$) $\chi_{V}(m_{i})\chi_{j}(m_{i})$, but this just says that the $i$-th column of $X(M)$ is left eigenvector of $N(V)$ with eigenvalue $\chi_{V}(m_{i})$. ∎
For $1\leq i\leq m$, let $C_{i,1},\dots,C_{i,k_{i}}$ denote the conjugacy classes of $G_{e_{i}}$, and pick a complete set of representatives $g_{i,1},\dots,g_{i,k_{i}}$ for these classes. Let $V$ be a $kM$-module with character $\chi_{V}$. By Lemma 1, when $p\nmid\lvert G_{e_{i}}\rvert,1\leq i\leq m$ each conjugacy class $C_{i,j}$ is canonically associated with a pair of left and right eigenvectors with eigenvalue $\chi_{V}(g_{i,j})$. We will refer to these as the $(i,j)$-th row of $(X(M))^{T}$ and the $(i,j)$-th column of $L^{-1}\overline{X}$, respectively. If $G_{e_{i}}$ is a maximal subgroup of $M$, extending notation from before we let $Z_{V}(G_{e_{i}})$ denote the elements $g\in G_{e_{i}}$ which act as scalars on $V$, and let $\omega_{V}(g)$ denote the corresponding scalar. We will write $k(n)$ for an asymptotic expression of $l(n)$.
Theorem 2.
@@
    (3) $l(n)=\sum_{i=1}^{m}\frac{1}{\lvert G_{e_{i}}\rvert}\sum_{j=1}^{k_{i}}\lvert C_{i,j}\rvert S_{i,j}\big(\chi(g_{i,j})\big)^{n}$
  2. 2.
@@
    (4) $l(n)\sim k(n):=\sum_{i=1}^{m}\frac{(\dim V)^{n}}{\lvert G_{e_{i}}\rvert}\sum_{j:g_{i,j}\in Z_{V}(G_{e_{i}})}\lvert C_{i,j}\rvert S_{i,j}\big(\omega_{V}(g_{i,j})\big)^{n}.$
  3. 3.
@@
For the eigenvalue $\chi(g_{i,j})$, choose the left eigenvector $w_{ij}^{T}$ to be the $(i,j)$-th row of $(X(M))^{T}$, which as recalled above is the same as $X^{T}L$, and choose the right eigenvector $v_{i,j}$ to be the $(i,j)$-th column of $L^{-1}\overline{X}$, normalized by a factor of $\lvert C_{i,j}\rvert/\lvert G_{e_{i}}\rvert$. From eigendecomposition we get
@@
Remark 4.
If $M=G$ is a group, then if $p\nmid\lvert G\rvert$ Theorem 2 gives
@@
where for $1\leq i\leq N$, $t_{i}$ is a representative for the $i$-th conjugacy class $C_{i}$ of $G$, and $S_{t}=\overline{S_{t}}$ is the sum over entries in column of the character table corresponding to $g_{t}$. Thus Theorem 2 generalizes the statement in [He25, Theorem 3]. $\Diamond$
Remark 5.
As we observed in the beginning of the section, if $kM$ is semisimple then $b(n)=l(n)$, so Theorem 2 gives a refinement of the result of Theorem 1 in this case. In particular, if no element other than $1$ act as identity on $M$, then the second statement of Theorem 2 recovers the formula of $a(n)$ in Theorem 1 in this case: this is because the submatrix of $L$ corresponding to the group of units $G$ is the identity matrix, so $S_{i,j}$ is exactly a column of the character table of $G$. Of course, if $kM$ is semisimple then we must have $p\nmid\lvert G_{e_{i}}\rvert$ for all $1\leq i\leq m$, so Theorem 2 applies (see e.g.[Ste16b, Theorem 5.19]). $\Diamond$